# version 0.75

### v4

@@ 12.3 Disturbance Response Controllers
The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC. These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time.
To simplify presentation, we only consider linear time invariant linear dynamical systems that are stable. This allows us to ignore the stabilizing controller component of disturbance action control and greatly simplifies the exposition of the main ideas. Extension to time-varying systems is explored in the exercises.
@@
A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by the matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t+1}=C\mathbf{x}^{\mathbf{nat}}_{t+1}.$
Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that are $\gamma$-stabilizing 111for example, by having additional linear controllers $K_{1:T}$ that are $\gamma$-stabilizing linear controllers for the system and
@@
If the perturbations are all zero, seemingly DRC policies produce a zero control. Therefore we could add a constant offset to the control, that can be incorporated as in Definition 6.5.
@@
We proceed to show that that the class of DRC approximates the class of LDC.
@@ 12.3.1 Expressivity of Disturbance Response Controllers
A DRC always produces controls that ensure the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
While DRCs are contained inside the policy class of DACs, they are still very expressive. Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
We have already seen that LDCs are approximately captured by GOACs in Lemma 12.5. It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of GOACs for linear time-invariant systems. Note that this observation in conjunction with Lemma 12.5 implies that DRCs also approximate LDCs.
To demonstrate that DRCs approximate GOACs, we instead prove that LOACs, a weaker class of policies, are $\varepsilon$-approximated by DRCs. Nevertheless, this observation is sufficient to conclude the stronger claim, given that GOACs can be written as LOACs on a lifted system as noted in Lemma 12.3, and simultaneously that a DRC with history $H^{\prime}$ on the lifted system can be written as a DRC with history $H^{\prime}+h$ over the original system.
@@
Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$. The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
@@
Recall that for the fully observed case the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations. It is this observation that allowed us to prove that DACs capture linear policies for fully observed systems. We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of LOACs can be written as linear functions of nature’s $y$’s. Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s (as DRCs do) allows us to encompass LOACs.
We introduce an augmented partially observed linear system that helps us write the observation-action pairs produced by a LOAC as a linear function of nature’s $y$’s. Crucially, the new system has the same nature’s $y$’s as the original system.
Lemma 12.9.
@@
It holds uniformly for all time $t$ that $\widetilde{\mathbf{y}^{\mathbf{nat}}_{t}}={\mathbf{y}^{\mathbf{nat}}_{t}}$ and further for any sequence of control inputs $\mathbf{u}_{t}$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a linear policy $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
@@
Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for $i>0$ and $M_{0}=K$, let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
@@
Given $\gamma$-stabilizability of the LOAC $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ suc that $\frac{\kappa}{\delta}\leq\gamma$ that
@@
$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left\|\sum_{j>1}CA^{j-1}B\mathbf{w}_{t-i-j}\right\|$
$\leq\frac{\kappa^{5}}{\delta}\sum_{i>h^{\prime}}\|(A+BKC)^{i-1}\|\leq\frac{\kappa^{6}\gamma}{\delta}(1-\frac{1}{\delta})^{h^{\prime}}$
@@
The approximation in terms of observations and controls implies $\varepsilon$-approximation of policy class for any Lipschitz cost sequence, concluding the proof. ∎
@@ 12.4 Summary
In this chapter, we extended our study of control to the partially observed setting, where the controller no longer has direct access to the system state but must instead rely on noisy or incomplete observations. This shift fundamentally changes the nature of the control problem, requiring the design of policies that can infer and act based on limited information.
We introduced a hierarchy of policy classes, starting with Linear Observation-Action Controllers (OACs), which directly map observations to actions, and progressing to Generalized LOACs (GOACs) and Linear Dynamic Controllers (LDCs), which incorporate history to improve decision-making. This increasing expressiveness highlighted a key limitation: memoryless policies are insufficient for optimal control, making structured state-dependent controllers necessary.
A major contribution of this chapter was the introduction of Disturbance Response Controllers (DRCs). Instead of explicitly estimating the system state, DRCs act directly on a counterfactual signal—the “Nature’s y’s”—capturing what the system would have observed in the absence of control inputs. We established that DRCs are a convex relaxation of LDCs, making them computationally attractive while maintaining strong expressive power. This result provides a principled approach to designing policies that are both efficient and robust in partially observed environments.
@@ 12.5 Bibliographic Remarks
The study of control in partially observed linear dynamical systems (PO-LDS) has a long history, with foundational contributions dating back to optimal estimation and control theories. This section provides an overview of key references related to the topics discussed in this chapter, including classical works, expressivity of policy classes, convex relaxations, and connections to online learning.
The problem of control under partial observation was first rigorously formulated in the context of the Linear Quadratic Gaussian (LQG) regulator, which combines Kalman filtering with optimal control [AM79]. The certainty equivalence principle, which asserts that optimal controllers can be designed assuming a perfect state estimate, was shown to have limitations in more general partially observed settings [Won68].
This chapter introduced a hierarchy of policy classes, including Linear Observation-Action Controllers (OACs), Generalized OACs (GOACs), Linear Dynamic Controllers (LDCs), and Disturbance Response Controllers (DRCs). The concept of disturbance response controllers was inspired by parameterizations used in robust control, particularly the Youla-Kucera parameterization [YJB76].
Convex relaxations have been a powerful tool for simplifying control design in PO-LDS. We have surveyed convex relaxations for learning in dynamical systems in the previous part of this book. Recently, spectral methods for control [AHL+18] have provided an alternative approach for convex relaxation, leveraging spectral filtering to bypass direct system identification.
The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC for control of partially observed systems. It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76]. For time-varying dynamics, the power of various policy classes was contrasted in [MGSH21].
@@ 12.6 Exercises
such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time. Additionally, verify that there exists such a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d. as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d. across time steps.
Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$. Also consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
@@
12.6. Construct a partially observed linear dynamical system subject to Gaussian i.i.d. perturbations with quadratic costs such that simultaneously it is true that (a) there exists a GOAC policy that incurs zero (or at most constant) aggregate cost, (b) yet any LOAC policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
@@
In the previous chapter, we introduced Disturbance Response Controllers (DRCs) as a structured class of policies for controlling partially observed linear dynamical systems (PO-LDS). Unlike classical observer-based controllers, which rely on explicit state estimation, DRCs act directly on Nature’s y’s, the system’s counterfactual response in the absence of control. While DRCs offer a powerful alternative to state-based policies, they did not yet incorporate online learning, leaving open the question of how to adapt them dynamically in uncertain environments.
@@
Unlike previous approaches that relied on fixed policy classes, GRC provides an adaptive control mechanism, bridging the gap between structured controllers and learning-based methods. A key focus of this chapter is extending policy regret minimization from the fully observed case to partial observability, introducing new challenges and solutions in learning-based control.
@@ 13.1 The Gradient Response Controller
Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter. The Gradient Response Controller algorithm is described in figure (8), for linear time-invariant dynamical systems with partial observation. Similar to the GPC, the main idea is to learn the parametrization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent. We henceforth prove that this parametrization is convex, allowing us to prove the main regret bound.
In algorithm (8) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising by playing DRC policy ${M}_{1:h}$ from the beginning of time, and same for $\mathbf{u}_{t}({M}_{1:h})$. Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 12.7, as
@@
6:  end for
@@
The GRC algorithm comes with a finite time performance guarantee: namely it guarantees vanishing worst case regret vs. the best disturbance response control policy in hindsight. As per the relationship between disturbance response control and other policies studied in section 12.3.1, this implies vanishing regret vs. linear dynamical controllers. The formal statement is given as follows.
@@
    The costs $c_{t}$ are convex, bounded and have bounded gradients w.r.t. both arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
@@
    The diameter of the constrained set for the parameters $M_{0:h}$ is bounded by $D$.
@@
    The matrices $\{{A},{B},C\}$ have bounded $\ell_{2}$ norm.
@@
As per Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded as a sublinear function of the number of iterations. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following conditions hold true:
@@
  • We need to ensure that optimizing the loss function $\ell_{t}$ as a function of the parameters $M_{0:h}$ is similar to optimizing the cost $c_{t}$ as a function of the observation and control.
First, we show that the loss function is convex w.r.t. to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
@@
Since the cost $c_{t}$ is a convex function w.r.t. its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 12.7,
@@
which is a linear function of the variables. Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and hence, the loss function is convex in ${M}_{0:h}$. ∎
Next, we use the following Lemma,
@@
13.1.1 Proof of Lemma 13.3
Proof.

### v5

@@ 12.3 Disturbance Response Controllers
The most important policy class in the context of nonstochastic control is that of Disturbance Response Controllers, or DRC. These policies act on a signal called Nature’s y’s, denoted $\mathbf{y}^{\mathbf{nat}}_{t}$, which is the would-be observation at time $t$ assuming that all inputs to the system were zero from the beginning of time.
For ease of exposition, we only consider linear time invariant linear dynamical systems that are intrinsically stable. This allows us to ignore the stabilizing controller component of disturbance response control and greatly simplifies the presentation of the main ideas. The extension to time-varying systems is explored in the exercises.
@@
A disturbance response policy for stable systems $\pi_{M_{1:h}}$ is parameterized by matrices $M_{0:h}=[{M}_{0},\ldots,{M}_{h}]$. It outputs control $\mathbf{u}_{t}^{\pi}$ at time $t$ according to the rule
@@
$\mathbf{x}^{\mathbf{nat}}_{t+1}$ $=A\mathbf{x}^{\mathbf{nat}}_{t}+\mathbf{w}_{t}=\sum_{i=0}^{t}A^{i}\mathbf{w}_{t-i}\ ,\ \mathbf{y}^{\mathbf{nat}}_{t}=C\mathbf{x}^{\mathbf{nat}}_{t}.$
Denote by $\Pi^{R}_{h,\gamma}$ the set of all disturbance response policies that satisfy [^1] Recall that for this section, we restrict our discussion to stable systems. Otherwise, we would also ask the policy class to be $\gamma$-stabilizing, which can, for example, be achieved by superimposing a fixed $\gamma$-stabilizing linear dynamical controller.
@@
If the perturbations are all zero, seemingly DRC policies produce zero control. Therefore, we could add a constant offset to the control that can be incorporated as in Definition 6.5.
@@
We proceed to show that the class of DRC approximates the class of LDC.
@@ 12.3.1 Expressivity of Disturbance Response Controllers
A DRC always produces controls that ensure that the system remains state-bounded. This can be shown analogously to Lemma 6.6 (see Exercise 12.6).
Although DRCs are contained within the policy class of DACs, they are still very expressive. Note that a DAC policy can not be executed in a partially observed system from observations alone; even the knowledge of the entire observation sequence $\mathbf{y}_{t}$ may not uniquely determine $\mathbf{w}_{t}$.
We have already seen that LDCs are approximately captured by generalized linear controllers in Lemma 12.5. It can be shown that the class of DRCs with appropriate parameters $\varepsilon$-approximates the class of generalized linear policies for linear time-invariant systems. This observation in conjunction with Lemma 12.5 implies that DRCs also approximate LDCs.
To demonstrate that DRCs approximate generalized linear controllers, we instead prove DRCs approximate linear controllers, a weaker class of policies. However, this is sufficient to conclude the stronger claim. To see this, first recall that generalized linear controllers acting on the last $h$ states can be written as linear controllers on a lifted system with $h$ states, as noted in Lemma 12.3. Together with the next lemma, this implies that a DRC on the lifted system can approximate generalized linear controllers on the original system. However, this is not quite what we promised. To wrap up the argument, one can verify that a DRC with history $h^{\prime}$ on the lifted system can be written as a DRC with history $h^{\prime}+h$ on the original system.
@@
Consider a $\gamma$-stable time-invariant linear dynamical system $(A,B,C)$. The class $\Pi^{R}_{h^{\prime},\gamma^{\prime}}$ $\varepsilon$-approximates the class $\Pi^{L}_{\gamma}$ for the parameters $h^{\prime}=O\left(\gamma\log\frac{\gamma}{\varepsilon}\right)$, $\gamma^{\prime}=O(\gamma^{6})$.
Recall that for fully observed LDS, the state-action pairs produced by execution of a linear policy can be written as linear functions of the perturbations. This observation allowed us to prove that DACs capture linear policies for fully observed systems. We prove an analogue for partially observed systems: The observation-action pairs produced by the execution of linear policies can be written as linear functions of nature’s $y$’s. Thereafter, we argue that parameterizing the controllers to be linear in nature’s $y$’s allows us to encompass linear observation-action policies.
@@
We introduce an augmented partially-observed linear system that helps us write the observation-action pairs produced by a linear policy as a linear function of nature’s $y$’s. In fact, we will shortly see that, for any sequence of control inputs, the augmented linear system produces the same sequence of obeservations as the original system. Thus, the new system has the same nature’s $y$’s as the original system.
@@
Fix any sequence of control inputs $\mathbf{u}_{t}$. Let $\mathbf{x}_{t},\mathbf{y}_{t}$ be the states and observations produced by the original partially-observed system. The equivalents for the augmented system are $[{\mathbf{x}^{\mathbf{nat}}_{t}}^{\top},\Delta_{t}^{\top}]^{\top},\tilde{\mathbf{y}}_{t}$, as we have described above. Now we can verify for all time $t$ that $\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}$ and $\mathbf{x}_{t}=\mathbf{x}^{\mathbf{nat}}_{t}+\Delta_{t}$.
Since this new system has the same input-output behavior as the original system, we might equivalently investigate how the execution of a linear controller $\pi$ affects the augmented system. Substituting $\mathbf{u}_{t}^{\pi}=K\mathbf{y}^{\pi}_{t}=KC(\mathbf{x}^{\mathbf{nat}}_{t}+\Delta^{\pi}_{t})$, the state-action sequence produced by a linear policy $\mathbf{u}_{t}^{\pi}=K\mathbf{y}_{t}^{\pi}$ can be written as
@@
Now, define $M_{i}=KC(A+BKC)^{i-1}BK$ for all positive $i\in[h^{\prime}]$ and $M_{0}=K$. These parameters specify our DRC. Let $\mathbf{y}_{t},\mathbf{u}_{t}$ be the observation-action sequence produced by this DRC. We have
@@
Given $\gamma$-stabilizability of the linear controller $K$ and $\gamma$-intrinsic stability of the of the system $(A,B,C)$, we have that for some $\kappa,\delta$ such that $\frac{\kappa}{\delta}\leq\gamma$,
@@
$\leq\sum_{i>h^{\prime}}\|KC(A+BKC)^{i-1}BK\|\left(\sum_{j>1}\left\|CA^{j-1}\mathbf{w}_{t-i-j}\right\|\right)$
$\leq\sum_{i>h^{\prime}}\|K\|^{2}\kappa(1-\delta)^{i-1}\sum_{j>1}\kappa(1-\delta)^{j}$
$\leq\frac{\kappa^{4}}{\delta}\sum_{i>h^{\prime}}(1-\delta)^{i-1}\leq\frac{\kappa^{4}}{\delta^{2}}(1-{\delta})^{h^{\prime}}\leq\frac{\kappa^{4}}{\delta^{2}}e^{-\delta h^{\prime}}$
@@
The approximation in terms of observations and controls implies $\varepsilon$-approximation of the policy class for any Lipschitz cost sequence, concluding the proof. ∎
@@ 12.4 Summary
In this chapter, we extended our study of control to the partially observed setting, where the controller no longer has direct access to the system state but must instead rely on noisy or incomplete observations. This shift fundamentally changes the nature of the control problem, requiring the design of policies that can infer and act on the basis of limited information.
We introduced a hierarchy of policy classes, starting with Linear Observation-Action Controllers, which directly map observations to actions, and progressing to generalized linear policies and Linear Dynamic Controllers (LDCs), which incorporate history to improve decision-making. This increasing expressiveness highlighted a key limitation: memoryless policies are insufficient for optimal control, making structured state-dependent controllers necessary.
A major contribution of this chapter was the introduction of Disturbance Response Controllers (DRCs). Instead of explicitly estimating the state of the system, DRCs act directly on a counterfactual signal, the “Nature’s y’s”, capturing what the system would have observed in the absence of control inputs. We established that DRCs are a convex relaxation of LDCs, making them computationally attractive while maintaining strong expressive power. This result provides a principled approach to designing policies that are efficient and robust in partially observed environments.
@@ 12.5 Bibliographic Remarks
The study of control in partially observed linear dynamical systems has a long history, with foundational contributions dating back to optimal estimation and control theories. This section provides an overview of key references related to the topics discussed in this chapter, including classical works, expressivity of policy classes, convex relaxations, and connections to online learning.
The problem of control under partial observation was first rigorously formulated in the context of the Linear Quadratic Gaussian (LQG) regulator, which combines Kalman filtering with optimal control [AM79]. The certainty equivalence principle, which states that optimal controllers can be designed assuming a perfect state estimate, was shown to have limitations in more general partially-observed settings [Won68].
This chapter introduced a hierarchy of policy classes, including linear observation-action controllers, generalized linear controllers, Linear Dynamic Controllers (LDCs), and Disturbance Response Controllers (DRCs). The concept of disturbance response controllers was inspired by parameterizations used in robust control, particularly the Youla-Kucera parameterization [YJB76].
Convex relaxations have been a powerful tool for simplifying control design in partially observed LDS. We have surveyed convex relaxations for learning in dynamical systems in the previous part of this book. Recently, spectral methods for control [AHL+18] have provided an alternative approach to convex relaxation, utilizing spectral filtering to bypass direct system identification.
The Disturbance Response Parametrization was introduced in [SSH20] as an analogue of DAC to control partially-observed systems. It can be viewed as a succinct state-space representation of the Youla parameterization [YJB76]. For time-varying dynamics, the power of various policy classes was contrasted in [MGSH21].
@@ 12.6 Exercises
such that for any sequence of $\mathbf{u}_{t}$, $\mathbf{y}_{t}=\mathbf{y}^{\prime}_{t}$ holds uniformly over time. Additionally, verify that there exists a construction of $\mathbf{z}_{t}$ where $\mathbf{w}^{\prime}_{t}$ are i.i.d. as long as $(\mathbf{w}_{t},\mathbf{v}_{t})$ are i.i.d. across time steps.
Hint: Consider the choice $\mathbf{w}^{\prime}_{t}=D^{\prime}_{t}\begin{bmatrix}\mathbf{w}_{t}\\ \mathbf{v}_{t}\end{bmatrix}$ for some $D^{\prime}_{t}$. Also, consider using a hidden state $\mathbf{z}_{t}$ with a larger number of coordinates than $\mathbf{x}_{t}$.
@@
12.6. Construct a partially observed linear dynamical system subject to Gaussian i.i.d. perturbations with quadratic costs such that simultaneously it is true that (a) there exists a genearlized linear policy that incurs zero (or at most constant) aggregate cost, (b) yet any linear observation-action policy incurs $\Omega(T)$ aggregate cost when run for $T$ time steps.
@@
In the previous chapter, we introduced Disturbance Response Controllers (DRCs) as a structured class of policies for controlling partially observed linear dynamical systems (PO-LDS). Unlike classical observer-based controllers, which rely on explicit state estimation, DRCs act directly on Nature’s y’s, the system’s counterfactual response in the absence of control. Although DRCs offer a powerful alternative to state-based policies, we have not yet integrated them into an online learning algorithm, leaving open the question of how to adapt them dynamically in uncertain environments.
@@
Unlike previous approaches that relied on fixed policy classes, GRC provides an adaptive control mechanism that bridges the gap between structured controllers and learning-based methods. A key focus of this chapter is to extend policy regret minimization from the fully observed case to partial observability, introducing new challenges and solutions in learning-based control.
@@ 13.1 The Gradient Response Controller
Partial observation requires modification to the Gradient Perturbation Controller we have studied in the previous chapter. The Gradient Response Controller algorithm is described in figure (8), for linear time-invariant dynamical systems with partial observation. Similarly to GPC, the main idea is to learn the parameterization of a disturbance response control policy using known online convex optimization algorithms, such as online gradient descent. We henceforth prove that this parameterization is convex, allowing us to prove the main regret bound.
In algorithm (8) we denote $\mathbf{y}_{t}({M}_{1:h})$ as the observation arising from playing the DRC policy ${M}_{1:h}$ from the beginning of time, and the same for $\mathbf{u}_{t}({M}_{1:h})$. Although it may be impossible to determine $\mathbf{w}_{t}$, especially if $C$ has fewer rows than columns, from the observations $\mathbf{y}_{t}$, $\mathbf{y}_{t}({M}_{1:h})$ can be determined from $\mathbf{y}^{\mathbf{nat}}_{t}$ alone by the application of Lemma 12.7, as
@@
5:     Construct loss $\ell_{t}({M}_{0:h})=c_{t}(\mathbf{y}_{t}({M}_{0:h}),\mathbf{u}_{t}({M}_{0:h}))$
6:     Update $M_{0:h}^{t+1}\leftarrow\prod_{\mathcal{K}}\left[{M}_{0:h}^{t}-\eta\nabla\ell_{t}({M}_{0:h}^{t})\right]$
7:  end for
@@
The GRC algorithm comes with a finite-time performance guarantee: it guarantees vanishing worst-case regret versus the best disturbance response control policy in hindsight. Following the relationship between disturbance response control and other policies studied in Section 12.3.1, this implies vanishing regret versus linear dynamical controllers. The formal statement is given as follows.
@@
    The costs $c_{t}$ are convex, bounded and have bounded gradients with respect to the arguments $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$.
@@
    The diameter of the constraint set for the parameters $M_{0:h}$ is bounded by $D$.
@@
    The matrices $\{{A},{B},C\}$ have bounded $\ell_{2}$ norms.
@@
According to Theorem 6.1, the online gradient descent algorithm for convex loss functions ensures that the regret is bounded as a sublinear function of the number of iterations. However, to directly apply the OGD algorithm to our setting, we need to make sure that the following conditions hold true:
@@
  • We must ensure that optimizing the loss function $\ell_{t}$ as a function of parameters $M_{0:h}$ is similar to optimizing the cost $c_{t}$ as a function of observation and control.
First, we show that the loss function is convex with respect to the variables ${M}_{0:h}$. This follows since the states and the controls are linear transformations of the variables.
@@
Since the cost $c_{t}$ is a convex function with respect to its arguments, we simply need to show that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ depend linearly on ${M}_{0:h}$. Using Lemma 12.7,
@@
which is a linear function of the variables. Thus, we have shown that $\mathbf{y}_{t}$ and $\mathbf{u}_{t}$ are linear transformations in ${M}_{0:h}$ and, hence, the loss function is convex in ${M}_{0:h}$. ∎
Next, we use the following lemma.
@@
13.1.1 Finishing up the Regret Bound
Proof of Lemma 13.3.
We first note that by the choice of step size by the Online Gradient Descent algorithm 3, as per Theorem 6.1, we have $\eta=\frac{D}{G\sqrt{T}}$. Thus,